OpenStudy (flexastexas):

2. A car accelerates uniformly from +10.0 m/s to +40.0 m/s over a distance of 125 m. How long did it take to go that distance

OpenStudy (flexastexas):

What formula do I use?

OpenStudy (aaronq):

use two of the kinematic formulas to eliminate acceleration (i.e. substitute one into the other), like when solving two simultaneous equations.

OpenStudy (flexastexas):

How would I know which to use?

OpenStudy (aaronq):

it doesn't matter which ones you use, as long as you use the information given \(v=v_0+at\rightarrow a=\dfrac{(v-v_0)}{t}\) then plug this into: \(v^2=v^2_0+2a\Delta x\) \(v^2=v^2_0+\dfrac{2(v-v_0)}{t}\Delta x\) solve for t

OpenStudy (flexastexas):

What do I plug into what? My initial velocity is 10m/s and final is 40m/s over a distance of 125m

OpenStudy (flexastexas):

@aaronq

OpenStudy (aaronq):

is this the first time you've seen these equations? \(v_0\) is the initial velocity \(v\) is the final velocity \(t\) is time \(\Delta x\) is the displacement

OpenStudy (thephysicsman):

acceleration is the change in velocity. it's the change in velocity over the change in time \[\frac{ v-v_0 }{ \Delta~t } = a \]